OpenStudy (anonymous):

what is the value of -29

OpenStudy (anonymous):

What is this supposed to mean?

OpenStudy (anonymous):

-29 = -29...

OpenStudy (k_lynn):

YOu mean like the absolute value? like this |29|?

OpenStudy (anonymous):

the question is what is the value of -29 a.29 b.29

OpenStudy (unanimoose):

It's just . . . -29 . . .

OpenStudy (k_lynn):

well if it doesn't say absolute value (|-29|) then -29=-29. But if it has these | lines in front and behind the number than |-29|= 29. its the same way with other numbers like -48. |-48|=48. 48=48.

OpenStudy (igreen):

Can you take a printscreen of your question? @haleigh101101

OpenStudy (unanimoose):

You're probably not able to paste part of the question.

OpenStudy (just_one_last_goodbye):

Since -29 = -29 then the answer would be A. -29

OpenStudy (unanimoose):

But both answers are just 29.

OpenStudy (unanimoose):

Options*

OpenStudy (just_one_last_goodbye):

she probably forgot to put the negative sign ^-^

OpenStudy (anonymous):

A.29 or B.-29

OpenStudy (just_one_last_goodbye):

so its B ^-^ -29

OpenStudy (unanimoose):

Then it's B. :P -29 = -29

OpenStudy (anonymous):

ok sorry i suck at math

OpenStudy (igreen):

The answer sounds too obvious..

OpenStudy (igreen):

I bet it's wrong.

OpenStudy (just_one_last_goodbye):

It ok lol xD we all have typo days ^-^ and honestly im horrible at math im surprised i even have a title "human calculator"

OpenStudy (anonymous):

can you help with this one 272+r -78 = ?

OpenStudy (just_one_last_goodbye):

simplify the following: r - 78 + 272

OpenStudy (just_one_last_goodbye):

Grouping like terms, r - 78 + 272 = r + (272 + 78): r + (272 - 78)

OpenStudy (anonymous):

I knew my answer was correct >;P

OpenStudy (just_one_last_goodbye):

272 - 78 = 194: \[Answer: r + 194\]